OpenStudy (anonymous):

Probability Question:

OpenStudy (anonymous):

Two identical biased coins are tossed together, and the outcome is recorded. After a large number of trials it is observed that the probability that both coins land showing tails is 0.49. What is the probability that both coins land showing heads?

OpenStudy (anonymous):

2/4 or 1/2

OpenStudy (anonymous):

not that easy (:

OpenStudy (anonymous):

@ParthKohli can try :p

OpenStudy (anonymous):

do you know the answer @Omniscience

OpenStudy (anonymous):

Obv.

Parth (parthkohli):

Experimental probability. I see.

OpenStudy (anonymous):

Try it (:

OpenStudy (anonymous):

why you are asking then @Omniscience

Parth (parthkohli):

An approach to this question is P(Heads) = 1 - P(Tails)

OpenStudy (anonymous):

Yes, That was what I thought but then that is not the first step.

Parth (parthkohli):

P(Heads) = 1 - 0.49

OpenStudy (anonymous):

lol, i did the same thing but it is not right.

Parth (parthkohli):

Is this experimental or theoretical probability?

Parth (parthkohli):

I go for experimental, as it says 'biased'

OpenStudy (anonymous):

lol, simple probability; very easy infact (:

Parth (parthkohli):

Is it somewhat like this? P(Coin will show a tails in one toss) = \(\sqrt{0.49}\)?

OpenStudy (anonymous):

Very close.

Parth (parthkohli):

Glad to hear that. Will the first step show the computation of the probability that the coin will show a 'tails' in one toss?

OpenStudy (anonymous):

Well, show the probability of throwing heads first.

Parth (parthkohli):

1/2 is theoretical. 1 - 0.49 is experimental.

Parth (parthkohli):

Well no... 1 - sqrt0.49

OpenStudy (anonymous):

Correct, which is the probability of throwing heads.

OpenStudy (anonymous):

Then what is the probabilty that both coins land showing heads?

OpenStudy (anonymous):

Wait 1-sqrt0.49; was the probability of head.

OpenStudy (anonymous):

one head*

Parth (parthkohli):

(1 - sqrt0.49)^2

Parth (parthkohli):

\(\Rightarrow 1^2 - 2(\sqrt{0.49})(1) + 0.49\)

OpenStudy (anonymous):

lol what

Parth (parthkohli):

\(\Rightarrow 1.49 - 2\sqrt{0.49} \)?

OpenStudy (anonymous):

you are over thinking.

Parth (parthkohli):

Omni, that is the identity lol \((a - b)^2 = a^2 - 2ab + b^2\)

Parth (parthkohli):

Idk... what else can I do here lol

OpenStudy (anonymous):

Don't overcomplicate the question!!

OpenStudy (anonymous):

Well, you got it right first; but you over complicated it. The answer is \(0.09\)

Parth (parthkohli):

But wait, that is the probability of a head in one toss.

OpenStudy (anonymous):

No that is the probability of two heads. Since once head is \(0.3\)

Parth (parthkohli):

Yep

OpenStudy (anonymous):

Do you get it?

Parth (parthkohli):

Yeah. 0.3 * 0.3

OpenStudy (anonymous):

yes; yaay! ^_^

Parth (parthkohli):

lol